# Emotions Where Art Thou: Understanding and Characterizing the Emotional Latent Space of Large Language Models

### v1

@@ 4.2 Universality Analysis
The first step of the universality analysis was to collect the mean-pooled hidden-states of the model when text $x_{i}$ from dataset $D$ is input into it. Then the mean-pooled hidden state is either projected onto the space described in Section 3 or statistics are directly derived from comparing the mean-pooled hidden states.
@@
Model Avg Cosine Similarity Avg MSE Avg Spectral Flatness Avg Frob Norm
Llama3.1-8B-Base 0.84 ± 0.12 1.79 ± 1.97 2.09 ± 0.39 7.70 ± 2.67
Llama3.1-8B-Instruct 0.93 ± 0.06 0.91 ± 1.07 2.29 ± 0.42 8.68 ± 6.36
Olmov2-Base 0.86 ± 0.15 1.87 ± 5.34 2.28 ± 0.40 8.02 ± 1.62
Olmov2-Instruct 0.90 ± 0.10 1.00 ± 2.17 2.26 ± 0.40 7.98 ± 1.31
Ministral 0.92 ± 0.04 1.44 ± 1.47 2.22 ± 0.44 7.60 ± 0.91
Table 2: Layer-wise average cosine similarity and linear alignment statistics by model.
Table 2 reports cosine similarity, regression error, and transformation statistics, each computed at every sublayer and then averaged across the full depth of the model. All models achieve a strong average centroid cosine similarity between emotions in real datasets and their synthetic counterparts (0.84–0.93 on average), showing that emotional categories are encoded in similar directions across domains and languages. Instruction-tuned models, especially LLaMA-3.1-Instruct and OLMo-v2-Instruct, show higher similarity and lower MSE than their base versions, indicating that tuning improves how closely real-emotion spaces align with the synthetic manifold. By contrast, spectral flatness and Frobenius norm values are similar across models and do not reliably separate base from instruct variants: LLaMA shows a slight increase with tuning, OLMo shows none or a decrease, and Ministral lies in the same range. This suggests that the improved alignment is not due to stronger or more isotropic transformations, but rather to better-aligned underlying representations.
@@
With the alignment between emotional spaces established, the distortion scores for each space of each model are assessed next. Results are summarized in Table  1. While traditional usage defines Stress-1 below 0.2 as acceptable and below 0.1 as strong Kruskal (1964), these thresholds were developed for low-dimensional embeddings (e.g., 2D). In the setting described here—mapping between high-dimensional hidden states—no canonical thresholds exist, but consistently low scores indicate robust relational preservation. This is most clearly observed in LLaMA-3.1-8B-Base, which achieves the lowest stress values across datasets. LLaMA-3.1-8B-Instruct and Ministral also yield relatively low Stress-1 and Stress-2 scores, though somewhat higher than the LLaMA-3.1-Base model, suggesting that instruction-tuned variants preserve emotional relations slightly less consistently. Interestingly, OLMo-v2 shows the opposite trend: the base model achieves lower stress than the instruction-tuned variant, but both versions exhibit substantially higher stress values than LLaMA or Ministral. This pattern suggests that OLMo encodes a less unified emotional space than the other models. Projection into the 50D synthetic subspace increases stress across all models, consistent with expected compression effects. For models with already low stress values, the increases remain modest and the geometry remains coherent, whereas for models with elevated stress (e.g., OLMo), dimensionality reduction does not mitigate distortion.
@@
To complement the stress scores, three high-dimensional embedding distortion metrics are reported: average distortion, $\ell_{2}$-distortion, and $\sigma$-distortion. Values near 1 for the first two indicate that pairwise distances are preserved up to global scaling, while low $\sigma$-distortion (ideal value =0=0) reflects stable proportionality in relative distances. These metrics highlight the contrast between base and instruction-tuned models. Both LLaMA-3.1-8B-Base and OLMo-v2-Base achieve distortion scores near ideal across most datasets (recall that no Ministral-Base model exists). For LLaMA-3.1-8B-Base, only two datasets deviate substantially from the ideal; excluding these outliers, its distortion metrics are more favorable than those of OLMo-v2-Base in both the full hidden-state space and the 50D synthetic subspace. See Appendix C for a contextualization of these scores.
@@
The instruction-tuned models, however, diverge considerably from their base counterparts. LLaMA-3.1-8B-Instruct produces roughly twice as many outlier datasets as LLaMA-3.1-8B-Base, while Ministral exhibits five “outliers” out of the nine datasets tested. Even after excluding these outliers, the distortion metrics for both models remain less favorable than those of their base equivalents in both full-space and 50D-space analyses, though still reasonably close to the ideal range. Nonetheless, for LLaMA-3.1-8B-Instruct and Ministral, the emotional embedding space remains partially universal: emotions are still represented in a broadly consistent manner across multiple datasets and languages. However, OLMo-v2-Instruct performs markedly worse. The spaces formed by emotion activations across different datasets are essentially mutually unintelligible. Its average distortion scores are extremely high across the entire model, suggesting that each dataset encodes emotions in a dataset- and language-specific fashion that resists mapping into a shared space.
@@
Linear probes trained on activations projected into the synthetic emotional latent space predict the emotions of human-written text at above-chance levels. These results indicate that emotional structure in LLMs is not only geometrically consistent but also linearly decodable across diverse domains, though probe performance varies with model family and the coherence of the underlying emotional space.
Comparing Table 2 with Table 1 underscores a central tension. The layer-averaged cosine similarities and regression errors suggest that all models align well with the synthetic emotional manifold, often with small differences across families. Yet the stress and distortion metrics reveal that, within the same models, relational structure can still be substantially warped—sometimes across large fractions of layers. This discrepancy reflects the fact that centroidal and regression-based measures capture global alignment, whereas stress and distortion expose finer-grained deviations in how relative distances between emotions are preserved. Thus, high apparent alignment at the aggregate level can coexist with local irregularities in the geometry of emotional spaces; however, linear probing shows that these spaces can still usefully and predictably predict emotion, even if there is some distortion between the synthetic and human-written spaces.
5 Model Psychology
The first perspective examined is neural encoding patterns. Using ML-AURA (Section 3), the selectively of individual neurons responses to emotional inputs is tested, treating each as a 1-vs-all classifier evaluated by AUROC. The percentage of neurons per layer exceeding $0.9$ AUROC are reported, taken as evidence of emotion-selective activation. Results focus on Llama3.1-8B-Base, with replications across models in Appendix D. Across the six Ekman emotions, an average of 75% of neurons per layer surpass this threshold, with sadness (98%) and surprise (97%) most pervasive, and fear lower (48%). This reflects sparse specialization rather than weak separability. Non-Ekman emotions—envy, neutral, excitement—also show strong separability, averaging 88%.
When analyzing by architectural component, MLP layers show slightly higher selectivity than attention layers ($79\%$ vs. $76.5\%$). Differentiability fluctuates across depth, with no clear monotonic trend: while the first layer starts at $76\%$ and the final layer ends at $76.3\%$, several peaks and troughs occur in between, with the highest selectivity observed at layer 26 ($79\%$). These patterns support the conclusion that emotional information is not confined to late layers or specialized regions, but is instead distributed broadly and redundantly throughout the network. These patterns are visualized by emotion in a layer-by-layer fashion in Figure 1.
To understand how emotions are geometrically represented in the network, we examine their structure within the derived SVD subspace. This subspace provides a low-dimensional lens into the model’s internal affective organization. Our first goal is to assess how consistently emotions are arranged along the principal axes across layers and layer types. To this end, we analyze the rank ordering of emotion centroids along each principal component, controlling for possible polarity flips.
Even when using a more fine-grained labeling scheme, as in the Go-Emotions dataset, which contains nearly three times as many emotion categories, we observe similar consistency. Rank-order correlations for Go-Emotions along the top three PCs remain high: Spearman values of $0.92$, $0.74$, and $0.73$, and Kendall’s Tau of $0.86$, $0.68$, and $0.68$. These findings reinforce the conclusion that the model’s emotional manifold is structurally stable, with interpretable axes. Appendix E reproduces the high consistency in how emotions are arranged along principal axes across all models studied.
Figure 2: Emotion centroids plotted on the emotional axis found
Having established the stability of the SVD subspace across layers and datasets, the semantic content of the leading principal components is examined. By analyzing the relative positions of emotion centroids along each axis, the underlying affective dimensions implicitly encoded by the model is inferred. Figure 2 visualizes the first three emotion axes that we describe below.
Figure 3: Plot of projected emotions in t-SNE space.
Figure 3 visualizes how emotional states are organized in hidden-state space using a 2D t-SNE projection of mean-pooled hidden states labeled with their respective emotions. Despite the dimensionality reduction, emotion classes form distinguishable, partially overlapping clusters, with closely related emotions (e.g., happy and excitement) frequently co-localized and others (e.g., fear and joy) appearing more spatially distant. While not all boundaries are sharp, the observed structure reinforces earlier findings: emotional information is embedded in a distributed yet semantically coherent geometry.
6 Steerability and the Limits of Control
Prior work on emotional steering in LLMs focuses primarily on shifting the emotional tone of generated text. Subramani et al. (2022) and Zou et al. (2023) learn vectors to modify output valence or categorical emotion. More recently, Hollinsworth et al. (2024) attempts to steer internal emotional representations but collapses emotion into a binary positive/negative axis, achieving valence shifts 53.5% of the time. In contrast, we aim for fine-grained control over the model’s internal perception of emotion across a full categorical space, while preserving semantic content.
Semantic Preservation.
Emotion Control.
Table 3 shows that the learned steering approach achieves consistent and accurate control over internal emotional representations across a diverse set of languages and datasets for all three models. For most emotions, post-steering accuracy typically exceeds 85% across models. Performance is robust even in multilingual settings, with particularly strong results in French, German, and Italian. Steerability remains high for many emotions in Hindi—a lower-resource language—suggesting that lexical sparsity and data imbalance remain limiting factors for certain emotions in under-resourced settings across models. The semantic similarity loss across all datasets, models, and emotions is low, indicating that steering preserves much of the original representational structure while enabling control of emotional perception. LLaMA-3.1-8B is the most steerable model for a plurality of emotions and datasets; however, Olmov2 shows the greatest average delta across emotions and datasets—an outcome notable given that Olmov2-Instruct was previously shown in Section 4 to struggle with representing emotions in a unified manner. Appendix F reports ablations isolating the method’s key factors. Finally, Table 4 illustrates how emotional steering manifests in practice, with inputs rewritten to express the target emotion.
Llama3.1-8B
Ministral
Olmov2
Dataset Original Text Shift Emotion Steering Rewrite
Table 4: Examples of text rewritten using the steering method.
7 Conclusion
These findings offer a structured account of how LLMs represent and modulate emotion. Future work should extend this analysis to multimodal models, investigating whether shared affective subspaces emerge across language, vision, and speech, and whether emotional representations in one modality can steer or constrain perception in another. Such models may yield a richer, more disentangled affective geometry, enabling both deeper interpretability and more naturalistic emotional reasoning.
Appendix A Stylistic Variations Among Selected Datasets
  1. 1.
    Go-Emotions Demszky et al. (2020): English Reddit comments
  2. 2.
    CARER Saravia et al. (2018): English tweets
  3. 3.
    SemEval-2007 Task 14 Strapparava & Mihalcea (2007): English news headlines
  4. 4.
    EmoEvent Plaza-del-Arco et al. (2020): English and Spanish tweets
  5. 5.
    Emotions in Drama Dennerlein et al. (2023): German plays from the 18th–19th century
  6. 6.
    Bhaav Kumar et al. (2019): Hindi short stories
  7. 7.
    MultiEmotions-It Sprugnoli et al. (2020): Italian YouTube and Facebook comments
  8. 8.
    EmoTextToKids Étienne et al. (2024): French journalistic and encyclopedic text aimed at children
Additionally, the GSM-8K dataset Cobbe et al. (2021) is used as a “negative relief” to provide context for the stress and distortion figures computed later on. This provides a reference point for interpreting stress and distortion scores in the absence of emotional alignment.
Table 5 summarizes the lexicographic and stylistic metrics for all selected datasets, except Bhaav, which is written in a low-resource language with limited tooling support for extracting such features. The table shows that the datasets contain a great diversity in style and complexity. The synthetic dataset has the longest sentences, highest syllables per word, and highest Flesch–Kincaid (FK) Grade Level Kincaid et al. (1975) score, making it among the most complex datasets to read. It is also the second most complex dataset in terms of the Dale–Chall readability score, which accounts for both average sentence length and the percentage of difficult words not on a list of familiar vocabulary Dale & Chall (1948). However, its type-token ratio (TTR) is the second lowest of all datasets, suggesting that despite its syntactic complexity, the vocabulary used is relatively constrained. Its intra-passage sentence length variability is relatively high, as indicated by a high sentence length standard deviation (sent len std), reflecting a mix of short and long constructions within the same passage.
Appendix B Cosine Similarities Between Datasets
Appendix C Distortion Score Contextualization
Appendix D ML-AURA Across Models

### v2

@@ 4.2 Universality Analysis
Using the hidden-state representations, projections, and metrics defined in Section 3, we evaluate how consistently emotions are encoded across datasets. Cosine similarity operates on emotion centroids, the regression MSE is derived from a linear map fitted to all mean-pooled-level activations sharing the same emotion label, and the stress/distortion metrics operate on cross-dataset, same-emotion distance pairs defined in the synthetic and real latent spaces.
Model Language Avg Cos Sim $\uparrow$ Stress-2 $\downarrow$ Avg Dist $\downarrow$ Probe Acc. $\uparrow$ Avg MSE $\downarrow$ Avg Spectral Flatness Avg Frob Norm
Llama-Base English 0.84 ± 0.13 0.15 ± 0.14 0.97 ± 0.22 0.47 ± 0.14 1.81 ± 1.99 2.04 ± 0.39 7.54 ± 1.82
Llama-Base Non-English 0.84 ± 0.12 0.18 ± 0.15 0.96 ± 0.22 0.4 ± 0.11 1.81 ± 2.00 2.10 ± 0.41 7.66 ± 2.45
Llama-Instruct English 0.93 ± 0.08 0.22 ± 0.11 0.78 ± 0.12 0.4 ± 0.06 0.93 ± 1.09 2.26 ± 0.41 8.70 ± 6.73
Llama-Instruct Non-English 0.94 ± 0.05 0.22 ± 0.15 1.01 ± 0.15 0.45 ± 0.06 0.89 ± 1.05 2.30 ± 0.41 8.66 ± 6.05
Olmov2-Base English 0.88 ± 0.13 0.59 ± 0.85 1.46 ± 0.63 0.42 ± 0.05 1.90 ± 5.40 2.19 ± 0.39 7.48 ± 0.41
Olmov2-Base Non-English 0.83 ± 0.16 0.61 ± 1.5 1.35 ± 0.78 0.45 ± 0.05 1.86 ± 5.30 2.35 ± 0.39 8.45 ± 2.05
Olmov2-Instruct English 0.90 ± 0.10 0.32 ± 0.32 47%* 0.47 ± 0.06 1.03 ± 2.24 2.20 ± 0.37 7.60 ± 0.64
Olmov2-Instruct Non-English 0.89 ± 0.09 0.43 ± 0.59 51%* 0.45 ± 0.05 0.97 ± 2.11 2.32 ± 0.40 8.30 ± 1.61
Ministral English 0.94 ± 0.06 0.21 ± 0.27 1.11 ± 0.17 0.39 ± 0.05 1.73 ± 2.29 2.17 ± 0.43 7.53 ± 0.73
Ministral Non-English 0.93 ± 0.09 0.24 ± 0.29 1.18 ± 0.12 0.45 ± 0.05 1.69 ± 2.25 2.24 ± 0.43 7.50 ± 0.83
Table 1: Per-model stress, distortion, linear-alignment metrics, and probe accuracy for English and non-English datasets. Lower distortion reflects greater geometric consistency. Cells marked with * indicate high stress/distortion; in these cases the table reports the percentage of layers exhibiting high distortion rather than raw scores. A full per-dataset breakdown appears in Appendix E.
Table 1 shows that all models exhibit high centroid cosine similarity between real and synthetic emotion directions (0.83–0.93), indicating stable cross-dataset encoding of emotional categories. English datasets are only marginally higher than non-English ones, suggesting near-equivalent representational fidelity across languages. Instruction-tuned variants show higher cosine similarity and lower regression error than their base models, reflecting closer alignment to the synthetic manifold. MSE values are broadly comparable across languages, with differences well within variance. By contrast, spectral flatness and Frobenius norms remain broadly similar across models and do not distinguish base from instruct variants: LLaMA shows a minor increase with tuning, OLMo shows none or a decrease, and Ministral remains in the same range. This pattern suggests that tuning improves representational alignment rather than relying on larger or more isotropic transformations.
Across datasets, LLaMA-3.1-8B-Base exhibits the lowest stress-2 values, indicating the most coherent emotional geometry (Table 1). Appendix E shows the same comparison broken down per dataset and includes additional stress and distortion metrics. Non-English datasets tend to have slightly higher stress-2 than their English counterparts. Average distortion varies on a model-by-model basis as to whether the English or non-English datasets have a higher score. LLaMA-3.1-8B-Instruct and Ministral also maintain relatively low Stress-2 scores, though both show a modest increase relative to the base model. OLMo-v2 displays a markedly different pattern: the base version shows substantially higher stress than LLaMA or Ministral, and the instruct variant is higher still, suggesting a less unified emotional space. Projecting representations into the 50D synthetic subspace increases stress for all models (shown in Appendix E), but the effect remains small for models with coherent geometry and is most pronounced for OLMo, where elevated stress persists after projection.
The average distortion metric further differentiates the models. LLaMA-3.1-8B-Base remains closest to the expected range for both English and non-English datasets, whereas OLMov2-Base shows substantially larger deviations. Distortion also behaves differently across languages: some models compress English representations (e.g., LLaMA-Instruct), others expand non-English representations (e.g., Ministral), and the direction of deviation is not consistent across base/instruct variants. For most models these shifts are modest, with OLMov2 exhibiting the most pronounced deviations. Across models, instruction-tuned variants generally show higher distortion than their base counterparts, with the notable exception of LLaMA-3.1-Instruct on non-English datasets, where distortion is nearly ideal. LLaMA-3.1-8B-Instruct therefore under-compresses English while improving non-English distortion, Ministral shows mild expansion while remaining stable, and OLMov2-Base has the largest distortion of any model, with OLMo-v2-Instruct pushing this even further, nearly half of its layers becoming severely over-distorted.
In terms of stress, however, the instruct variant of OLMo exhibits markedly lower stress than its base model. Ministral’s stress-2 score is lower still, aligning closely with LLaMA-Instruct. The comparatively low stress-2 of OLMo-v2-Instruct alongside its extreme distortion suggests that instruction tuning improves global directional alignment while simultaneously degrading local geometric coherence, indicating that OLMo’s emotional manifold may be more fragile and easily warped or over-rotated by tuning. Appendix E provides dataset-level breakdowns and additional stress and distortion measures, with further contextualization in Appendix D.
Certain dataset–model combinations exhibit extremely high distortion across large portions of the network when evaluated relative to the synthetic emotional space. These patterns are detailed in Appendix E. Even in cases with elevated distortion, however, substantial subsets of layers continue to preserve a coherent and transferable emotional geometry.
Linear probes achieve above-chance accuracy when predicting the emotions of human-written text from projected activations (Table 1) . This indicates that emotional structure in LLMs is not only geometrically consistent but also linearly decodable across diverse domains, though probe performance varies with model family and the coherence of the underlying emotional space.
Comparing the alignment metrics in Table 1 with the geometric faithfulness and structural stability metrics underscores a central tension. The layer-averaged cosine similarities and regression errors suggest that all models align well with the synthetic emotional manifold, often with small differences across families. Yet the stress and distortion metrics reveal that, within the same models, relational structure can still be substantially warped—sometimes across large fractions of layers. This discrepancy reflects the fact that centroidal and regression-based measures capture global alignment, whereas stress and distortion expose finer-grained deviations in how relative distances between emotions are preserved. Thus, high apparent alignment at the aggregate level can coexist with local irregularities in the geometry of emotional spaces; however, linear probing shows that these spaces can still usefully and predictably predict emotion, even if there is some distortion between the synthetic and human-written spaces. Taken together, these results indicate that the claims concern the directional organization of the emotional subspace rather than strict isometry, and that the local warping revealed by stress and distortion is expected in high-dimensional compression while remaining compatible with a functionally coherent and manipulable emotional manifold.
4.3 Model Psychology
The first perspective examined is neural encoding patterns using ML-AURA (Section 3). Results focus on Llama3.1-8B-Base, with replications across models in Appendix F. Across the six Ekman emotions, an average of 75% of neurons per layer exceed the selectivity threshold, with sadness (98%) and surprise (97%) most pervasive, and fear lower (48%). This reflects sparse specialization rather than weak separability, and, importantly, selectivity rather than activation magnitude: ML-AURA identifies neurons whose responses discriminate emotions, not neurons that merely fire strongly. Non-Ekman emotions—envy, neutral, excitement—also show strong separability, averaging 88%.
When analyzing by architectural component, MLP layers show slightly higher selectivity than attention layers ($79\%$ vs. $76.5\%$). Selectivity fluctuates across depth, with no clear monotonic trend: while the first layer starts at $76\%$ and the final layer ends at $76.3\%$, several peaks and troughs occur in between, with the highest selectivity observed at layer 26 ($79\%$). These patterns support the conclusion that emotional information is not confined to late layers or specialized regions, but is instead distributed broadly and redundantly throughout the network. These patterns are visualized by emotion in a layer-by-layer fashion in Figure 1.
To understand how emotions are geometrically represented in the network, we examine their structure within the derived SVD subspace. This subspace provides a low-dimensional lens into the model’s internal affective organization. Our first goal is to assess how consistently emotions are arranged along the principal axes across layers and layer types.
Figure 2: Emotion centroids plotted on the emotional axis found
Even when using a more fine-grained labeling scheme, as in the Go-Emotions dataset, which contains nearly three times as many emotion categories, we observe similar consistency. Rank-order correlations for Go-Emotions along the top three PCs remain high: Spearman values of $0.92$, $0.74$, and $0.73$, and Kendall’s Tau of $0.86$, $0.68$, and $0.68$. These findings reinforce the conclusion that the model’s emotional manifold is structurally stable, with interpretable axes. Appendix G reproduces the high consistency in how emotions are arranged along principal axes across all models studied.
Having established the stability of the SVD subspace across layers and datasets, the semantic content of the leading principal components is examined. Figure 2 visualizes the first three emotion axes that we describe below.
Figure 3: Plot of projected emotions in t-SNE space.
Figure 3 provides a t-SNE visualization of the emotional space. Despite the dimensionality reduction, emotion classes form distinguishable, partially overlapping clusters, with closely related emotions (e.g., happy and excitement) frequently co-localized and others (e.g., fear and joy) appearing more spatially distant. While not all boundaries are sharp, the observed structure reinforces earlier findings: emotional information is embedded in a distributed yet semantically coherent geometry.
4.4 Steering Emotion Representations
Prior work on emotional steering in LLMs focuses primarily on shifting the emotional tone of generated text. (Subramani et al., 2022) and (Zou et al., 2023) learn vectors to modify output valence or categorical emotion. More recently, (Hollinsworth et al., 2024) attempts to steer internal emotional representations but collapses emotion into a binary positive/negative axis, achieving valence shifts 53.5% of the time. In contrast, we aim for fine-grained control over the model’s internal perception of emotion across a full categorical space, while preserving semantic content.
Semantic Preservation.
Emotion Control.
The steering module is evaluated by measuring how reliably it shifts the model’s emotion classification toward the target label while preserving semantic similarity. Table 2 shows that the learned steering approach achieves consistent and accurate control of internal emotional representations across English and non-English datasets for all three models. For most emotions, post-steering accuracy exceeds 80%, with many cases reaching 90–100%. Even for harder emotions and lower-resource languages, steering reliably improves target classification, with nearly all settings achieving at least  50% accuracy. Semantic similarity losses remain low across models and languages, indicating that steering preserves the core representational content while shifting affective interpretation. LLaMA-3.1-8B is the most steerable model across a plurality of settings, while OLMo-v2 shows the largest average improvement—despite OLMo-v2-Instruct previously exhibiting substantial geometric misalignment in Section 4. Appendix H reports full results per dataset, and Appendix I provides ablations isolating key methodological factors. Table 3 illustrates how emotional steering manifests in practice, with inputs rewritten to express the target emotion.
Model Language Sad Happy Fear Anger Neutral Disgust Envy Excitement Surprise Overall
Llama3.1-8B English 22 $\rightarrow$ 99 (0.14) 20 $\rightarrow$ 92 (0.23) 7 $\rightarrow$ 68 (0.25) 21 $\rightarrow$ 58 (0.21) 1 $\rightarrow$ 96 (0.21) 0 $\rightarrow$ 88 (0.24) 0 $\rightarrow$ 92 (0.24) 0 $\rightarrow$ 90 (0.23) 8 $\rightarrow$ 64 (0.22) 9 $\rightarrow$ 83 (0.22)
Llama3.1-8B Non-English 9 $\rightarrow$ 100 (0.12) 13 $\rightarrow$ 88 (0.21) 5 $\rightarrow$ 68 (0.21) 14 $\rightarrow$ 65 (0.20) 16 $\rightarrow$ 95 (0.20) 3 $\rightarrow$ 94 (0.21) 0 $\rightarrow$ 99 (0.21) 1 $\rightarrow$ 88 (0.21) 7 $\rightarrow$ 64 (0.19) 8 $\rightarrow$ 84 (0.20)
Olmov2 English 27 $\rightarrow$ 93 (0.09) 32 $\rightarrow$ 98 (0.07) 2 $\rightarrow$ 75 (0.05) 21 $\rightarrow$ 100 (0.07) 6 $\rightarrow$ 100 (0.04) 1 $\rightarrow$ 99 (0.04) 0 $\rightarrow$ 91 (0.04) 0 $\rightarrow$ 81 (0.07) 7 $\rightarrow$ 83 (0.07) 11 $\rightarrow$ 91 (0.06)
Olmov2 Non-English 33 $\rightarrow$ 84 (0.08) 21 $\rightarrow$ 80 (0.07) 2 $\rightarrow$ 59 (0.04) 23 $\rightarrow$ 100 (0.06) 8 $\rightarrow$ 99 (0.05) 1 $\rightarrow$ 98 (0.04) 0 $\rightarrow$ 84 (0.04) 4 $\rightarrow$ 73 (0.08) 4 $\rightarrow$ 91 (0.07) 11 $\rightarrow$ 86 (0.06)
Ministral English 17 $\rightarrow$ 76 (0.09) 24 $\rightarrow$ 90 (0.10) 5 $\rightarrow$ 98 (0.09) 17 $\rightarrow$ 100 (0.10) 16 $\rightarrow$ 100 (0.10) 2 $\rightarrow$ 96 (0.10) 0 $\rightarrow$ 99 (0.10) 1 $\rightarrow$ 82 (0.10) 7 $\rightarrow$ 99 (0.10) 10 $\rightarrow$ 91 (0.10)
Ministral Non-English 14 $\rightarrow$ 47 (0.09) 28 $\rightarrow$ 60 (0.09) 5 $\rightarrow$ 90 (0.08) 22 $\rightarrow$ 97 (0.09) 16 $\rightarrow$ 95 (0.09) 3 $\rightarrow$ 87 (0.10) 0 $\rightarrow$ 90 (0.10) 3 $\rightarrow$ 51 (0.09) 4 $\rightarrow$ 88 (0.09) 11 $\rightarrow$ 79 (0.09)
Table 2: Top-1 prediction rates before and after learned steering for each target emotion across datasets and the semantic similarity loss between the hidden-state representations before and after steering. Each cell shows baseline $\rightarrow$ post-steering (average semantic similarity loss) accuracy.
Dataset Original Text Shift Emotion Steering Rewrite
Table 3: Examples of text rewritten using the steering method.
5 Conclusion
These findings offer a structured account of how LLMs represent and modulate emotion. Future work should extend this analysis to multimodal models, investigating whether shared affective subspaces emerge across language, vision, and speech, and whether emotional representations in one modality can steer or constrain perception in another. Such models may yield a richer, more disentangled affective geometry, enabling both deeper interpretability and more naturalistic emotional reasoning. Another interesting future direction is to examine the developmental trajectory of emotional representations during pretraining, although doing so would require access to intermediate training checkpoints of large-scale models.
Limitations. Our universality claim is conditioned on the language and style being reasonably well represented in the model’s pretraining data. In out-of-distribution settings, such as nineteenth-century German drama or low-resource Hindi, the emotional latent space shows somewhat higher distortion and stress, yet remains directionally coherent, with probe accuracy above chance and successful steering. These results indicate that even under limited pretraining coverage, the learned emotional geometry retains usable structure rather than collapsing. Importantly, the steering experiments in Appendix H show that usable control remains achievable even in these regimes: Hindi, the weakest-performing language, still supports an absolute steering shift of roughly 50% on average across models, while German theater texts exhibit average shifts exceeding 80% across models. These results indicate that, although representational fidelity degrades when pretraining coverage is sparse, the underlying emotional manifold does not collapse; instead, it retains sufficient structure to support both detection and targeted intervention.
Ethics Statement. Because our method enables controlled shifts in a model’s internal affective representation, we acknowledge the need to articulate its appropriate use and its limitations. The steering mechanism is intentionally constrained: it acts only on intermediate hidden states, preserves semantic content through an explicit alignment regularizer, and cannot induce or amplify harmful content beyond what the base model already permits. Nonetheless, certain emotions, especially high-arousal or interpersonal ones such as anger or contempt, may yield more forceful stylistic rewrites, reflecting asymmetries in how models encode these emotions. We therefore recommend applying steering only in transparent, user-directed contexts, such as tone adjustment, therapeutic or reflective writing tools, accessibility interfaces, or cross-emotion normalization for evaluation. The method is not intended for covert style manipulation, persuasion, or emotionally charged rewriting without user consent. Finally, variability in steerability across languages and domains (e.g., low-resource or archaic corpora) functions as an inherent boundary: the technique does not uniformly override model behavior but respects representational limits.
Appendix A Appendix
Appendix B Stylistic Variations Among Selected Datasets
  1. 1.
    Go-Emotions (Demszky et al., 2020): English Reddit comments
  2. 2.
    CARER (Saravia et al., 2018): English tweets
  3. 3.
    SemEval-2007 Task 14 (Strapparava and Mihalcea, 2007): English news headlines
  4. 4.
    EmoEvent (Plaza-del-Arco et al., 2020): English and Spanish tweets
  5. 5.
    Emotions in Drama (Dennerlein et al., 2023): German plays from the 18th–19th century
  6. 6.
    Bhaav (Kumar et al., 2019): Hindi short stories
  7. 7.
    MultiEmotions-It (Sprugnoli and others, 2020): Italian YouTube and Facebook comments
  8. 8.
    EmoTextToKids (Étienne et al., 2024): French journalistic and encyclopedic text aimed at children
Additionally, the GSM-8K dataset (Cobbe et al., 2021) is used as a “negative relief” to provide context for the stress and distortion figures computed later on. This provides a reference point for interpreting stress and distortion scores in the absence of emotional alignment.
Table 4 summarizes the lexicographic and stylistic metrics for all selected datasets, except Bhaav, which is written in a low-resource language with limited tooling support for extracting such features. The table shows that the datasets contain a great diversity in style and complexity. The synthetic dataset has the longest sentences, highest syllables per word, and highest Flesch–Kincaid (FK) Grade Level (Kincaid et al., 1975) score, making it among the most complex datasets to read. It is also the second most complex dataset in terms of the Dale–Chall readability score, which accounts for both average sentence length and the percentage of difficult words not on a list of familiar vocabulary (Dale and Chall, 1948). However, its type-token ratio (TTR) is the second lowest of all datasets, suggesting that despite its syntactic complexity, the vocabulary used is relatively constrained. Its intra-passage sentence length variability is relatively high, as indicated by a high sentence length standard deviation (sent len std), reflecting a mix of short and long constructions within the same passage.
Appendix C Cosine Similarities Between Datasets
Appendix D Distortion Score Contextualization
Appendix E Universality Expanded
@@
In Section 4, stress-2 and average distortion scores for the models are discussed, with results compared across English and non-English datasets. In this appendix, the corresponding per-dataset results for stress, distortion, and probe accuracy are presented, along with an expanded set of stress and distortion metrics.
@@
Table 5 presents the stress and distortion score per-model and per-dataset. LLaMA-3.1-8B-Base, which achieves the lowest stress values across datasets. LLaMA-3.1-8B-Instruct and Ministral also yield relatively low Stress-1 and Stress-2 scores, though somewhat higher than the LLaMA-3.1-Base model, suggesting that instruction-tuned variants preserve emotional relations slightly less consistently. Interestingly, OLMo-v2 shows the opposite trend: the base model achieves lower stress than the instruction-tuned variant, but both versions exhibit substantially higher stress values than LLaMA or Ministral. This pattern suggests that OLMo encodes a less unified emotional space than the other models. Projection into the 50D synthetic subspace increases stress across all models, consistent with expected compression effects. For models with already low stress values, the increases remain modest and the geometry remains coherent, whereas for models with elevated stress (e.g., OLMo), dimensionality reduction does not mitigate distortion.
@@
To complement the stress scores, three high-dimensional embedding distortion metrics are reported: average distortion, $\ell_{2}$-distortion, and $\sigma$-distortion. These metrics highlight the contrast between base and instruction-tuned models. Both LLaMA-3.1-8B-Base and OLMo-v2-Base achieve distortion scores near ideal across most datasets (recall that no Ministral-Base model exists). For LLaMA-3.1-8B-Base, only two datasets deviate substantially from the ideal; excluding these outliers, its distortion metrics are more favorable than those of OLMo-v2-Base in both the full hidden-state space and the 50D synthetic subspace. See Appendix D for a contextualization of these scores.
@@
The instruction-tuned models diverge considerably from their base counterparts. LLaMA-3.1-8B-Instruct produces roughly twice as many outlier datasets as LLaMA-3.1-8B-Base, while Ministral shows five outliers out of the nine datasets tested. Even after excluding these outliers, distortion remains higher for both models in both full-space and 50D-space analyses. Despite this, LLaMA-3.1-8B-Instruct and Ministral still maintain broadly consistent emotional structure across multiple datasets and languages. OLMov2-Instruct, however, performs markedly worse: the emotional spaces induced by different datasets are largely incompatible, with distortion elevated across nearly all layers and datasets.
@@
These per-dataset analyses show that while stress and distortion vary across model families and datasets, a broadly coherent emotional geometry is retained across datasets and languages, especially in the more stable base models. Probe accuracy trends reinforce this pattern, indicating that despite local geometric failures, large portions of each model preserve a shared, dataset-invariant emotional subspace.
Appendix F ML-AURA Across Models